# Mitigating Manipulation and Enhancing Persuasion: A Reflective Multi-Agent Approach for Legal Argument Generation

### v1

@@ Abstract.
Large Language Models (LLMs) are increasingly explored for legal argument generation, yet they pose significant risks of manipulation through hallucination and ungrounded persuasion, and often fail to utilize provided factual bases effectively or abstain when arguments are untenable. This paper introduces a novel reflective multi-agent method designed to address these challenges in the context of legally compliant persuasion. Our approach employs specialized agents—a Factor Analyst and an Argument Polisher—in an iterative refinement process to generate 3-ply legal arguments (plaintiff, defendant, rebuttal). We evaluate Reflective Multi-Agent against single-agent, enhanced-prompt single-agent, and non-reflective multi-agent baselines using four diverse LLMs (GPT-4o, GPT-4o-mini, Llama-4-Maverick-17b-128e, Llama-4-Scout-17b-16e) across three legal scenarios: “arguable”, “mismatched”, and “non-arguable”. Results demonstrate Reflective Multi-Agent’s significant superiority in successful abstention (preventing generation when arguments cannot be grounded), marked improvements in hallucination accuracy (reducing fabricated and misattributed factors), particularly in “non-arguable” scenarios, and enhanced factor utilization recall (improving the use of provided case facts). These findings suggest that structured reflection within a multi-agent framework offers a robust computable method for fostering ethical persuasion and mitigating manipulation in LLM-based legal argumentation systems, a critical step towards trustworthy AI in law. Project page: Link
LLM Evaluation, Legal Argument Generation, Ethical AI, Multi-Agent Systems, Reflection Mechanisms, Hallucination Mitigation, Abstention, Case-Based Reasoning
copyright: nonejournalyear: 2025doi: conference: Workshop on Legally Compliant Intelligent Chatbots at ICAIL 2025; June 16, 2025; Chicago, ILisbn: ccs: Computing methodologies Natural language processingccs: Applied computing Law
@@ 1. Introduction
Argumentation, at its core, is the art of persuasion, employing logical reasoning and evidence to influence an audience on a particular matter (nickerson2020argumentation). In the legal domain, argumentation takes on a specialized form, where the careful construction of evidence-based claims is crucial. The ability to persuade effectively, while adhering to ethical and factual standards, is fundamental to legal practice (rhode1985ethical). As LLMs become increasingly refined, their potential to assist in, and even automate, aspects of legal argument generation brings both immense opportunities and significant challenges, particularly concerning the integrity of persuasion in this high-stakes field.
@@ 1.1. The Rise of LLMs in Legal Domain
The integration of LLMs into the legal domain presents a paradigm shift, offering considerable potential for automating routine tasks, assisting in legal research, document drafting, and even argument generation (siino2025exploring; gray2025generating; zhang2025measuring). LLMs can swiftly sift through vast legal data, draft legal documents, and assist in complex legal analysis, thereby enabling lawyers to focus on more strategic aspects of their work (nielsen2024building). However, the inherent persuasive capabilities that make these models valuable also introduce a complex dichotomy: while they can be leveraged for beneficial applications, they simultaneously pose significant risks of manipulation and unethical influence if not carefully managed (bommasani2021opportunities). The primary challenge lies in harnessing their power for ethical assistance while mitigating the potential for misuse.
These risks are amplified by the fact that the persuasive influence of LLMs is not unidirectional. Beyond their role as persuaders, these systems can also be susceptible to persuasion themselves, rendering them vulnerable to adversarial attacks and the reinforcement of biases present in their training data or input prompts (bozdag2025readsystematicsurveycomputational). For instance, an LLM tasked with generating legal arguments might have been trained on datasets containing systemic biases or could be influenced by subtly crafted adversarial inputs during its operational lifecycle (draper2023potential). This “persuadee” vulnerability means that an LLM could inadvertently generate arguments that, while appearing coherent and persuasive, perpetuate these biases or reflect manipulated information, thereby subtly distorting legal outcomes. Such complex interplay between an LLM’s persuasive output and its susceptibility to influence necessitates robust internal validation mechanisms within LLM-based legal chatbots. These mechanisms must go beyond mere output generation to meticulously scrutinize the factual grounding, logical consistency, and potential biases of the arguments produced, ensuring their integrity and reliability.
@@ 1.2. The Triad of Challenges
Pilot studies and broader research have identified critical shortcomings in LLM performance when applied to legal domain including argument generation (dahl2024large; gray2025generating; zhang2025measuring). Three primary challenges emerge: hallucination, inadequate abstention, and poor factor utilization.
Hallucination, the generation of text that is factually incorrect, inconsistent, or not supported by input data, is particularly pernicious in the legal field where precision and accuracy are fundamental. Studies have reported alarmingly high hallucination rates by LLMs in response to legal queries, with models fabricating case details or misstating legal principles (magesh2024hallucination). Such inaccuracies can lead to nonsensical or harmful legal advice (krook2024large).
Abstention, the ability of a model to refrain from answering when an argument is ungroundable or when it lacks sufficient information, is crucial for reliability. However, LLMs often fail to recognize the boundaries of their knowledge or the untenability of a query, proceeding to generate responses even when they should abstain (wen2024know).
Factor utilization refers to the extent to which an LLM incorporates relevant factual elements (factors) from provided case materials into its generated arguments (nadler1983evidence). Poor factor utilization results in arguments that may be superficially plausible but lack substantive grounding in the specific facts of the case, thereby diminishing their persuasive strength and utility.
These three challenges are often interlinked. A failure to abstain in scenarios where an argument cannot be legitimately grounded (“mismatched” or “non-arguable” cases) significantly increases the likelihood of hallucination. If a model is compelled to generate an argument without a proper factual basis, it is more prone to inventing facts or misapplying existing ones to fulfill the generation task. This directly translates to a higher risk of manipulation, as the generated argument may appear coherent but be based on falsehoods or irrelevant information.
@@ 2.1. LLMs for Legal Reasoning and Argument Generation
The application of LLMs to legal tasks, including argument generation, is growing (siino2025exploring; gray2025generating; zhang2025measuring). Computational legal argument models, especially those using case-based reasoning and ‘factors’ (stereotypical fact patterns), offer a strong foundation for this endeavor, with systems like HYPO pioneering factor-based case analysis (ashley1990), followed by developments in factor hierarchies (CATO (aleven1997teaching)) and value incorporation (grabmair2017predicting). These factor-based methods are key for evaluating argument factuality, central to our study. However, LLMs face challenges in specialized legal argument construction (padiu2024extent) and are prone to “legal hallucinations”—generating factually incorrect or misapplied legal content (dahl2024hallucinating)—highlighting the need for domain-specific adaptations and robust verification, as addressed by our Reflective Multi-Agent framework, which aims to enhance the grounding and accuracy of LLM-generated legal arguments, building on prior work to enhance LLM performance in law.
@@ 2.2. Computational Models of Persuasion and Manipulation in AI
The persuasive capabilities of AI systems, particularly LLMs, are increasingly recognized, with models functioning as persuaders, being susceptible to persuasion themselves (persuadees), and even acting as arbiters of persuasive attempts (breum2024persuasive). This multifaceted role presents opportunities for beneficial applications but concurrently introduces significant risks of manipulation, social engineering, and unethical influence (hu2025j). A central challenge, especially within sensitive domains such as law, is the development of “computable methods for ensuring ethical persuasion” (rogiers2024persuasion). The concept of “computational manipulation” is particularly germane to legal AI. As explored in contexts like AI-driven persuasive technology in contract law, AI systems can exploit user data and cognitive biases to influence decisions, potentially undermining autonomous decision-making processes without explicit user awareness (faraoni2024contract; burtell2023artificial). This notion directly correlates with problematic LLM behaviors, such as the generation of “favorable” yet false factual assertions or the vigorous defense of an untenable position. Our work seeks to mitigate such manipulative tendencies by enhancing factual grounding and promoting abstention when arguments lack merit.
@@ 2.3. Multi-Agent Systems in Complex Reasoning and Dialogue
LLM-based Multi-Agent Systems (MAS) are emerging as a potent paradigm for addressing complex problems, leveraging distributed task handling and specialized agent roles to enhance robustness and reasoning capabilities (guo2024large; gao2024agentscope). Multi-agent debate frameworks, for instance, have been proposed to improve the quality of reasoning by fostering critical interaction among agents, thereby potentially counteracting issues like the “Degeneration-of-Thought” problem observed in single-agent systems (liang2023encouraging). The proposed Reflective Multi-Agent system builds upon these principles by employing distinct agents for Plaintiff and Defendant roles. Furthermore, it introduces specialized sub-agents—the Factor Analyst and Argument Polisher—which collaborate to meticulously refine arguments. This architecture represents a nuanced application of Multi-Agent Systems, designed to harness the benefits of distributed expertise and structured dialogue for improved legal argument generation.
@@ 2.4. Reflection, Self-Correction, and Iterative Refinement in LLMs
Mechanisms enabling reflection and self-correction are crucial for enhancing the reliability and accuracy of LLM outputs (pan2023automatically). However, the efficacy of unaided self-correction remains a subject of debate. The “SELF-[IN]CORRECT” hypothesis, for example, posits that LLMs may not be consistently better at discriminating the quality of their own responses than they are at generating initial ones (huang2023large). This suggests that structured and guided reflection mechanisms may prove more effective. Challenges inherent in self-correction, such as difficulties in error detection and inherent self-bias, further motivate the development of more robust and systematic approaches (xi2024enhancing). The Reflective Multi-Agent framework’s reflective process is designed to be explicit and role-based, employing specialized agents (the Factor Analyst and Argument Polisher) for distinct analytical and refinement tasks. This structured critique aims to overcome the limitations associated with unguided self-correction. This approach aligns with findings from frameworks like “Reflexion,” which demonstrate the power of structured reflection in improving LLM performance (shinn2023reflexion).
@@ 2.5. Metrics for Evaluating LLM Outputs
The rigorous evaluation of LLM-generated content, particularly in high-stakes domains like law, necessitates robust and domain-relevant metrics. For assessing hallucination, a critical concern, various benchmarks and detection methods are continuously being developed (ji2023survey; luo2024hallucination; ravi2024lynx). The distinction between extrinsic and intrinsic hallucination, as offered by frameworks like “HalluLens” (bang2025hallulens), is particularly relevant; our work primarily addresses intrinsic hallucination by ensuring arguments are grounded in provided case factors. The capacity for appropriate abstention—refraining from providing an answer when it cannot be adequately supported—is another critical aspect of LLM reliability (feng2024don). Surveys and studies provide comprehensive overviews of abstention methods and evaluation metrics, highlighting the importance of this capability (wen2024know; madhusudhan2024llms). Furthermore, evaluating factor utilization, or the faithfulness of generated arguments to source material, is essential for ensuring the substantive quality of legal outputs (liu2023trustworthy). Frameworks like ICAT, which evaluate factual accuracy and coverage by decomposing text into claims and aligning them with relevant aspects (factors), offer valuable methodologies (samarinas2025beyond). QA-based verification methods (fabbri2021qafacteval) share conceptual similarities with our approach of employing an external LLM for factor summarization and comparison. This aligns with the broader trend of using “LLM-as-a-Judge” approaches for nuanced evaluations, a field with its own set of evolving best practices and considerations (zheng2023judging; gu2024survey). Our evaluation methodology draws upon these established principles to provide an assessment of the Reflective Multi-Agent framework.
@@ 3.1. Architectural Overview
The Reflective Multi-Agent framework is designed to generate a 3-ply legal argument structure (ashley1990; zhang2025measuring):
@@ 3.2. Agent Roles
  • Functions: Detect Hallucinations (identify factors not in input cases c1, c2, c3), Mandate Abstention (output ”TERMINATE” if argument is untenable).
@@ 3.3. The Iterative Reflection Workflow
    Final Output for Ply: ”TERMINATE” if mandated by Factor Analyst. Otherwise, the polished initial argument (if no revision) or re-polished revised argument.
@@ 4.1. Task Definition
The core task is to generate a 3-ply legal argument. The input are three cases (c1: current case, c2: Plaintiff’s precedent, c3: Defendant’s precedent) represented by legal factors. These legal factors are derived from foundational work in legal AI focusing on U.S. trade secret misappropriation law (ashley1990). A standardized set of 26 factors is used, encapsulating key factual aspects of a case, such as circumstances surrounding disclosure, security measures implemented, characteristics of the information, and relevant employee conduct. Each factor is designated as typically favoring either the plaintiff (P) or the defendant (D). For instance, a case might be represented as:
@@
This structured factor representation provides the essential ground truth for subsequent automated evaluation by allowing objective comparison between cases based on shared and distinguishing factors. The process follows a structure established by (ashley1990). First, the system, acting as the Plaintiff, argues for a favorable outcome by citing one of the provided precedent cases (c2 or c3) as analogous, emphasizing shared pro-plaintiff factors between the current case (c1) and the cited precedent. Second, the system takes on the role of the Defendant. It responds by distinguishing the precedent case cited by the Plaintiff (highlighting differing factors) and then presents the other precedent case as a counter-example that favors the defendant, focusing on shared pro-defendant factors between c1 and this counter-example precedent. Third, the system acts again as the Plaintiff to deliver a rebuttal. This involves distinguishing the counter-example case cited by the Defendant and reinforcing the original argument by emphasizing factors that differentiate the current case from the Defendant’s chosen precedent. The expected output is either the structured Plaintiff’s Argument (based on c1 and c2), the Defendant’s Counterargument (based on c1 and c3), and the Plaintiff’s Rebuttal (based on c1 and c2), or a ”TERMINATE” signal if the system determines that a valid argument cannot be constructed based on the provided factors.
For the argument generation, a max_tokens limit of 1,000 was set, which proved sufficient for the typical length of a 3-ply argument structure. Other standard parameters included top_p=1, frequency_penalty=0, and presence_penalty=0. The factor extraction step, performed by the external evaluator LLM (GPT-4.1), also utilized fixed deterministic settings (temperature=0, top_p=1) to ensure consistency in the evaluation process itself.

### v2

@@ Abstract.
Large Language Models (LLMs) are increasingly explored for legal argument generation, yet they pose significant risks of manipulation through hallucination and ungrounded persuasion, and often fail to utilize provided factual bases effectively or abstain when arguments are untenable. This paper introduces a novel reflective multi-agent method designed to address these challenges in the context of legally compliant persuasion. Our approach employs specialized agents (factor analyst and argument polisher) in an iterative refinement process to generate 3-ply legal arguments (plaintiff, defendant, rebuttal). We evaluate reflective multi-agent against single-agent, enhanced-prompt single-agent, and non-reflective multi-agent baselines using four diverse LLMs (GPT-4o, GPT-4o-mini, Llama-4-Maverick-17b-128e, Llama-4-Scout-17b-16e) across three legal scenarios: “arguable”, “mismatched”, and “non-arguable”. Results demonstrate that the reflective multi-agent approach excels at successful abstention by preventing generation when arguments cannot be grounded, improves hallucination accuracy by reducing fabricated and misattributed factors and enhances factor utilization recall by better using the provided case facts. These findings suggest that structured reflection within a multi-agent framework offers a robust method for fostering ethical persuasion and mitigating manipulation in LLM-based legal argumentation systems.
Project Page: lizhang-aiandlaw.github.io/A-Reflective-Multi-Agent-Approach-for-Legal-Argument-Generation
Trustworthy AI, Multi-Agent Systems, Hallucination Mitigation, Abstention, Legal Argument Generation
copyright: nonejournalyear: 2025conference: 2nd ConventicLe on Artificial Intelligence Regulation and Safety at ICAIL 2025; June 16, 2025; Chicago, ILccs: Computing methodologies Natural language processingccs: Applied computing Law
@@ 1. Introduction
Argumentation, at its core, is the art of persuasion, employing logical reasoning and evidence to influence an audience on a particular matter (nickerson2020argumentation, 31). In the legal domain, argumentation takes on a specialized form, where the careful construction of evidence-based claims is crucial. The ability to persuade effectively, while adhering to ethical and factual standards, is fundamental to legal practice (rhode1985ethical, 36). As LLMs become increasingly refined, their potential to assist in, and even automate, aspects of legal argument generation brings both immense opportunities and challenges, particularly concerning the integrity of persuasion in this high-stakes field.
@@ 1.1. The Rise of LLMs in Legal Domain
The integration of LLMs into the legal domain presents a paradigm shift, offering considerable potential for automating routine tasks, assisting in legal research, document drafting, and even argument generation (siino2025exploring, 40, 17, 46). LLMs can swiftly sift through vast legal data, draft legal documents, and assist in complex legal analysis, thereby enabling lawyers to focus on more strategic aspects of their work (nielsen2024building, 32). However, the inherent persuasive capabilities that make these models valuable also introduce a complex dichotomy: while they can be leveraged for beneficial applications, they simultaneously pose risks of manipulation and unethical influence if not carefully managed (bommasani2021opportunities, 4). The primary challenge lies in harnessing their power for ethical assistance while mitigating the potential for misuse.
These risks are amplified by the fact that the persuasive influence of LLMs is not unidirectional. Beyond their role as persuaders, these systems can also be susceptible to persuasion themselves, rendering them vulnerable to adversarial attacks and the reinforcement of biases present in their training data or input prompts (bozdag2025readsystematicsurveycomputational, 5). For instance, an LLM tasked with generating legal arguments might have been trained on datasets containing systemic biases or could be influenced by subtly crafted adversarial inputs during its operational lifecycle (draper2023potential, 11). This “persuadee” vulnerability means that an LLM could inadvertently generate arguments that, while appearing coherent and persuasive, perpetuate these biases or reflect manipulated information, thereby subtly distorting legal outcomes. Such complex interplay between an LLM’s persuasive output and its susceptibility to influence necessitates robust internal validation mechanisms within LLM-based legal chatbots. These mechanisms must go beyond mere output generation to meticulously scrutinize the factual grounding, logical consistency, and potential biases of the arguments produced, ensuring their integrity and reliability.
@@ 1.2. The Triad of Challenges
Pilot studies and broader research have identified shortcomings in LLM performance when applied to legal domain including argument generation (dahl2024large, 9, 17, 46). Three challenges emerge: hallucination, inadequate abstention, and poor factor utilization.
Hallucination, the generation of text that is factually incorrect, inconsistent, or not supported by input data, is particularly pernicious in the legal field where precision and accuracy are fundamental. Studies have reported alarmingly high hallucination rates by LLMs in response to legal queries, with models fabricating case details or misstating legal principles (magesh2024hallucination, 29). Such inaccuracies can lead to nonsensical or harmful legal advice (krook2024large, 24).
Abstention, the ability of a model to refrain from answering when an argument is ungroundable or when it lacks sufficient information, is crucial for reliability. However, LLMs often fail to recognize the boundaries of their knowledge or the untenability of a query, proceeding to generate responses even when they should abstain (wen2024know, 43).
Factor utilization refers to the extent to which an LLM incorporates relevant factual elements (factors) from provided case materials into its generated arguments (nadler1983evidence, 30). Poor factor utilization results in arguments that may be superficially plausible but lack substantive grounding in the specific facts of the case, thereby diminishing their persuasive strength and utility.
These three challenges are often interlinked. A failure to abstain in scenarios where an argument cannot be legitimately grounded (“mismatched” or “non-arguable” cases) increases the likelihood of hallucination. If a model is compelled to generate an argument without a proper factual basis, it is more prone to inventing facts or misapplying existing ones to fulfill the generation task. This directly translates to a higher risk of manipulation, as the generated argument may appear coherent but be based on falsehoods or irrelevant information.
@@ 2.1. LLMs for Legal Reasoning and Argument Generation
The application of LLMs to legal tasks, including argument generation, is growing (siino2025exploring, 40, 17, 46). Computational legal argument models, especially those using case-based reasoning and ‘factors’ (stereotypical fact patterns), offer a strong foundation for this endeavor, with systems like HYPO pioneering factor-based case analysis (ashley1990, 2), followed by developments in factor hierarchies (CATO (aleven1997teaching, 1)) and value incorporation (grabmair2017predicting, 16). These factor-based methods are key for evaluating argument factuality, central to our study. However, LLMs face challenges in specialized legal argument construction (padiu2024extent, 33) and are prone to “legal hallucinations”—generating factually incorrect or misapplied legal content (dahl2024hallucinating, 8)—highlighting the need for domain-specific adaptations and robust verification, as addressed by our Reflective Multi-Agent framework, which aims to enhance the grounding and accuracy of LLM-generated legal arguments, building on prior work to enhance LLM performance in law.
@@ 2.2. Computational Models of Persuasion and Manipulation in AI
The persuasive capabilities of AI systems, particularly LLMs, are increasingly recognized, with models functioning as persuaders, being susceptible to persuasion themselves, and even acting as arbiters of persuasive attempts (breum2024persuasive, 6). This multifaceted role presents opportunities for beneficial applications but concurrently introduces risks of manipulation, social engineering, and unethical influence (hu2025j, 21). A central challenge, especially within sensitive domains such as law, is the development of “computable methods for ensuring ethical persuasion” (rogiers2024persuasion, 37). The concept of “computational manipulation” is particularly germane to legal AI. As explored in contexts like AI-driven persuasive technology in contract law, AI systems can exploit user data and cognitive biases to influence decisions, potentially undermining autonomous decision-making processes without explicit user awareness (faraoni2024contract, 13, 7). This notion directly correlates with problematic LLM behaviors, such as the generation of “favorable” yet false factual assertions or the vigorous defense of an untenable position. Our work seeks to mitigate such manipulative tendencies by enhancing factual grounding and promoting abstention when arguments lack merit.
@@ 2.3. Multi-Agent Systems in Complex Reasoning and Dialogue
LLM-based Multi-Agent Systems (MAS) are emerging as a potent paradigm for addressing complex problems, leveraging distributed task handling and specialized agent roles to enhance robustness and reasoning capabilities (guo2024large, 19, 15). Multi-agent debate frameworks, for instance, have been proposed to improve the quality of reasoning by fostering interaction among agents, thereby potentially counteracting issues like the “Degeneration-of-Thought” problem observed in single-agent systems (liang2023encouraging, 25). The proposed Reflective Multi-Agent system builds upon these principles by employing distinct agents for Plaintiff and Defendant roles. Furthermore, it introduces specialized sub-agents—the Factor Analyst and Argument Polisher—which collaborate to meticulously refine arguments. This architecture represents a nuanced application of Multi-Agent Systems, designed to harness the benefits of distributed expertise and structured dialogue for improved legal argument generation.
@@ 2.4. Reflection, Self-Correction, and Iterative Refinement in LLMs
Mechanisms enabling reflection and self-correction are crucial for enhancing the reliability and accuracy of LLM outputs (pan2023automatically, 34). However, the efficacy of unaided self-correction remains a subject of debate. The “SELF-[IN]CORRECT” hypothesis, for example, posits that LLMs may not be consistently better at discriminating the quality of their own responses than they are at generating initial ones (huang2023large, 22). This suggests that structured and guided reflection mechanisms may prove more effective. Challenges inherent in self-correction, such as difficulties in error detection and inherent self-bias, further motivate the development of more robust and systematic approaches (xi2024enhancing, 45). The Reflective Multi-Agent framework’s reflective process is designed to be explicit and role-based, employing specialized agents (the Factor Analyst and Argument Polisher) for distinct analytical and refinement tasks. This structured critique aims to overcome the limitations associated with unguided self-correction, which demonstrates the power of structured reflection in improving LLM performance (shinn2023reflexion, 39).
@@ 2.5. Metrics for Evaluating LLM Outputs
The rigorous evaluation of LLM-generated content, particularly in high-stakes domains like law, necessitates robust and domain-relevant metrics. For assessing hallucination, various benchmarks and detection methods are continuously being developed (ji2023survey, 23, 27, 35). The distinction between extrinsic and intrinsic hallucination, as offered by frameworks like “HalluLens” (bang2025hallulens, 3), is particularly relevant; our work primarily addresses intrinsic hallucination by ensuring arguments are grounded in provided case factors. The capacity for appropriate abstention is another aspect of LLM reliability (feng2024don, 14). Surveys and studies provide comprehensive overviews of abstention methods and evaluation metrics, highlighting the importance of this capability (wen2024know, 43, 28). Furthermore, evaluating factor utilization, or the faithfulness of generated arguments to source material, is essential for ensuring the substantive quality of legal outputs (liu2023trustworthy, 26). Frameworks like ICAT, which evaluate factual accuracy and coverage by decomposing text into claims and aligning them with relevant aspects (factors), offer valuable methodologies (samarinas2025beyond, 38). QA-based verification methods (fabbri2021qafacteval, 12) share conceptual similarities with our approach of employing an external LLM for factor summarization and comparison. This aligns with the broader trend of using “LLM-as-a-Judge” approaches for nuanced evaluations, a field with its own set of evolving best practices and considerations (zheng2023judging, 47, 18). Our evaluation methodology draws upon these established principles to provide an assessment of the Reflective Multi-Agent framework.
@@ 3.1. Architectural Overview
The Reflective Multi-Agent framework is designed to generate a 3-ply legal argument structure (ashley1990, 2, 46):
@@ 3.2. Agent Roles
  • Functions: Detect Hallucinations (identify factors not in input cases c1, c2, c3), Mandate Abstention (output “TERMINATE” if argument is untenable).
@@ 3.3. The Iterative Reflection Workflow
    Final Output for Ply: “TERMINATE” if mandated by Factor Analyst. Otherwise, the polished initial argument (if no revision) or re-polished revised argument.
@@ 4.1. Task Definition
The core task is to generate a 3-ply legal argument. The input are three cases (c1: current case, c2: Plaintiff’s precedent, c3: Defendant’s precedent) represented by legal factors. These legal factors are derived from foundational work in legal AI focusing on U.S. trade secret misappropriation law (ashley1990, 2). A standardized set of 26 factors is used, encapsulating key factual aspects of a case, such as circumstances surrounding disclosure, security measures implemented, characteristics of the information, and relevant employee conduct. Each factor is designated as typically favoring either the plaintiff (P) or the defendant (D). For instance, a case might be represented as:
@@
This structured factor representation provides the essential ground truth for subsequent automated evaluation by allowing objective comparison between cases based on shared and distinguishing factors. The process follows a structure established by (ashley1990, 2). First, the system, acting as the Plaintiff, argues for a favorable outcome by citing one of the provided precedent cases (c2 or c3) as analogous, emphasizing shared pro-plaintiff factors between the current case (c1) and the cited precedent. Second, the system takes on the role of the Defendant. It responds by distinguishing the precedent case cited by the Plaintiff (highlighting differing factors) and then presents the other precedent case as a counterexample that favors the defendant, focusing on shared pro-defendant factors between c1 and this counterexample precedent. Third, the system acts again as the Plaintiff to deliver a rebuttal. This involves distinguishing the counterexample case cited by the Defendant and reinforcing the original argument by emphasizing factors that differentiate the current case from the Defendant’s chosen precedent. The expected output is either the structured Plaintiff’s Argument (based on c1 and c2), the Defendant’s Counterargument (based on c1 and c3), and the Plaintiff’s Rebuttal (based on c1 and c2), or a “TERMINATE” signal if the system determines that a valid argument cannot be constructed based on the provided factors.
For the argument generation, a ‘‘max_tokens’’ limit of 1,000 was set, which proved sufficient for the typical length of a 3-ply argument structure. Other standard parameters included ‘‘top_p’’=1, ‘‘frequency_penalty’’=0, and ‘‘presence_penalty’’=0. The factor extraction step, performed by the external evaluator LLM (GPT-4.1), also utilized fixed deterministic settings (‘‘temperature’’=0, ‘‘top_p’’=1) to ensure consistency in the evaluation process itself.